# LADR: Locality-Aware Dynamic Rescue for Efficient Text-to-Image Generation with Diffusion Large Language Models

### 13 Mar 2026 v1

@@ 3.1 Preliminaries: Discrete Diffusion and Variational Bounds
Let $\mathbf{z}_{0}=[z_{1},\dots,z_{N}]\in\mathcal{V}^{N}$ represent a discrete image sequence flattened from a $H\times W$ feature map, where each token belongs to a codebook $\mathcal{V}$. The discrete diffusion process is a forward Markov chain $q(\mathbf{z}_{t}|\mathbf{z}_{t-1})$ that progressively corrupts $\mathbf{z}_{0}$ by replacing tokens with a special $\texttt{MASK}]$ token. The marginal distribution at time $t\in[0,1]$ is given by:
@@
$\text{where }q(z_{t,i}=\texttt{[M]}|z_{0,i})=\gamma(t),$ (1)
@@
$\mathcal{L}\approx\mathbb{E}_{t,\mathbf{z}_{0}}\left[-\sum_{i\in\mathcal{M}_{t}}\log p_{\theta}(z_{0,i}|\mathbf{z}_{\mathcal{O}_{t}})\right],$ (2)
where $\mathcal{O}_{t}$ denotes the set of observed indices. During inference, iterative decoding approximates the joint distribution via conditional independence assumption: $p_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})\approx\prod_{i\in\mathcal{M}_{t}}p_{\theta}(z_{i}|\mathbf{z}_{\mathcal{O}_{t}})$. Standard acceleration methods strictly follow $\gamma(t)$, discarding potentially correct predictions in early stages.
@@ 3.2.3 Manifold Consistency via Inverse Scheduling
Remark. This necessitates our Inverse Scheduling technique, which acts as a temporal projection operator to correct the trajectory after aggressive rescue operations.
@@ 4.2 Main Results
Setting Avg. t (s) $\downarrow$ Global Entity Attribute Relation Other Overall $\uparrow$
@@
Beyond efficiency, our results indicate that the significant reduction in sampling steps does not come at the cost of visual fidelity, while it yields a highly competitive performance profile across diverse benchmarks. Conversely, we note a performance drop in fine-grained high-frequency tasks, such as the text rendering score in UniGen-Bench, suggesting that the model remains sensitive to the reduction of iterative refinement steps in some scenarios. Overall, the experiment results demonstrate that our method achieves a superior efficiency-fidelity trade-off, delivering efficient decoding speeds while preserving robust generative capabilities in image generation.
@@ 4.3 Ablation Studies and Analysis
Figure 5 presents the quantitative comparison on the GenEval benchmark. We observe that the Non-Neighbor strategy yields the lowest performance (Overall 84.05), significantly lagging behind our method (87.30). This confirms that forcing the model to resolve isolated tokens early—even those with high confidence gaps—leads to error propagation, as these predictions may lack sufficient spatial grounding. Interestingly, the random strategy (86.49) outperforms the Non-Neighbor variant but it is still lagging behind the proposed accelerated method. This indicates that our Spatial-Aware strategy provides the optimal balance, ensuring that the accelerated decoding trajectory respects the structural dependencies of the image.
@@ 5 Conclusion and Future Work
While LADR demonstrates the potential of exploiting image spatial locality for acceleration of parallel decoding, our current method still has some limitations. The implementation relies on empirically determined hyperparameters, such as the confidence threshold $\tau$ and rescue ratios $\alpha$. These values were selected to validate the core hypothesis that spatial neighbors facilitate faster convergence rather than to locate the global optimum.
@@ A.1 Proof of Theorem 1 (Margin-based Error Bound)
We aim to find the maximum possible error, which corresponds to minimizing the confidence $p_{(1)}$ subject to the constraints.
@@ 1. Define Constraints: The probability distribution must sum to 1, and the margin condition must hold:
$\sum_{k=1}^{K}p_{(k)}$ $=1,$ (12)
$p_{(1)}-p_{(2)}$ $\geq\tau.$ (13)
2. Worst-Case Analysis: To maximize error ($1-p_{(1)}$), we must minimize $p_{(1)}$. This occurs when the remaining probability mass $1-p_{(1)}$ is concentrated as much as possible in the “competitor” classes to satisfy the constraints tightest.
From Eq. (13), we express $p_{(2)}$ in terms of $p_{(1)}$:
$p_{(2)}\leq p_{(1)}-\tau.$ (14)
3. Bounding the Sum: Since $p_{(1)}$ is the maximum, all other probabilities $p_{(k)}$ for $k>2$ are also bounded by $p_{(2)}$ (and thus $p_{(1)}-\tau$), but a stricter bound for the binary case (or focusing on the top-2 dominating case) gives the tightest constraint on $p_{(1)}$.
Consider the relation between the top-2 probabilities and the total sum:
$p_{(1)}+p_{(2)}$ $\leq\sum_{k=1}^{K}p_{(k)}=1.$ (15)
4. Deriving the Inequality: Substitute the margin constraint $p_{(2)}\leq p_{(1)}-\tau$ into the sum inequality:
$p_{(1)}+(p_{(1)}-\tau)$ $\leq p_{(1)}+p_{(2)}\leq 1$ (16)
$2p_{(1)}-\tau$ $\leq 1$ (17)
$2p_{(1)}$ $\leq 1+\tau$ (18)
$p_{(1)}$ $\leq\frac{1+\tau}{2}.$ (19)
Correction for Lower Bound of Correctness: Wait, we are looking for the minimum valid $p_{(1)}$ to find the maximum error. The constraint is $p_{(1)}+p_{(2)}+\sum_{k=3}^{K}p_{(k)}=1$. To minimize $p_{(1)}$, we need to maximize the other terms. The largest possible value for $p_{(2)}$ is $p_{(1)}-\tau$.
Thus, in the worst case (where probability is split maximally between top-2):
$p_{(1)}+(p_{(1)}-\tau)$ $\approx 1\quad(\text{assuming }\sum_{k=3}^{K}p_{(k)}\approx 0)$ (20)
$2p_{(1)}$ $\approx 1+\tau$ (21)
@@
Any $p_{(1)}$ smaller than $\frac{1+\tau}{2}$ would require $p_{(2)}>p_{(1)}-\tau$ to sum to 1, violating the margin condition.
Therefore, the lower bound on confidence is:
$p_{(1)}\geq\frac{1+\tau}{2}.$ (23)
5. Calculating the Error Bound: The probability of error is the complement of the top-1 probability:
$P(\mathcal{E})$ $=1-p_{(1)}$ (24)
$\leq 1-\frac{1+\tau}{2}$ (25)

### 10 Apr 2026 v2

@@ 3.1 Preliminaries: Discrete Diffusion and Variational Bounds
Let $\mathbf{z}_{0}=[z_{0,1},\dots,z_{0,N}]\in\mathcal{V}^{N}$ represent a discrete image sequence flattened from a $H\times W$ feature map, where each token belongs to a codebook $\mathcal{V}$. The discrete diffusion process is a forward Markov chain $q(\mathbf{z}_{t}|\mathbf{z}_{t-1})$ that progressively corrupts $\mathbf{z}_{0}$ by replacing tokens with a special $\texttt{MASK}]$ token. The marginal distribution at time $t\in[0,1]$ is given by:
@@
$\text{where }q(z_{t,i}=\texttt{[MASK]}|z_{0,i})=\gamma(t),$ (1)
@@
$\mathcal{L}\approx\mathbb{E}_{t,\mathbf{z}_{0}}\left[-\sum_{i\in\mathcal{M}_{t}}\log p_{\theta}(z_{0,i}|\mathbf{z}_{t,\mathcal{O}_{t}})\right],$ (2)
where $\mathcal{O}_{t}$ denotes the set of observed indices. During inference, iterative decoding approximates the joint distribution via conditional independence assumption: $p_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})\approx\prod_{i\in\mathcal{M}_{t}}p_{\theta}(z_{0,i}|\mathbf{z}_{t,\mathcal{O}_{t}})$. Standard acceleration methods strictly follow $\gamma(t)$, discarding potentially correct predictions in early stages.
@@ 3.2.3 Manifold Consistency via Inverse Scheduling
Remark. This necessitates our Inverse Scheduling technique, which acts as a temporal projection operator to correct the trajectory after aggressive rescue operations. $\gamma^{-1}$ is the inverse function of $\gamma$.
@@ 4.2 Main Results
Method Avg. t (s) $\downarrow$ Global Entity Attribute Relation Other Overall $\uparrow$
@@
Beyond efficiency, our results indicate that the significant reduction in sampling steps does not come at the cost of visual fidelity, while it yields a highly competitive performance profile across diverse benchmarks. Conversely, we note a performance drop in fine-grained high-frequency tasks, such as the text rendering score in UniGen-Bench, suggesting that the model remains sensitive to the reduction of iterative refinement steps in some scenarios. To further investigate this, we conducted an additional experiment on UniGen-Bench’s text category exploring a delayed rescue strategy. Specifically, we completely disabled the rescue operation during the Exploration Phase ($t_{eff}<0.2$) and kept LADR’s other settings unchanged. As illustrated in Table 5, the text rendering score recovers to 26.10, which is comparable to the base model’s 27.87. Meanwhile, the inference time is 22.51s, still maintaining a highly efficient 2.5x speedup. This demonstrates that LADR could achieve a highly adaptable efficiency-quality trade-off. Overall, the experiment results demonstrate that our method achieves a superior efficiency-fidelity trade-off, delivering efficient decoding speeds while preserving robust generative capabilities in image generation.
Method Avg.t (s) $\downarrow$ Text
Lumina-DiMOO 57.21 27.87
LADR 13.94 16.38
LADR(delayed rescue) 22.51 26.10
Table 5: Efficiency and performance comparison of different strategies on the text category of UniGen-Bench.
@@ 4.3 Ablation Studies and Analysis
Figure 5 presents the quantitative comparison on the GenEval benchmark. We observe that the Non-Neighbor strategy yields the lowest performance (Overall score: 84.05), significantly lagging behind our method, which achieves an overall score of 87.30. This confirms that forcing the model to resolve isolated tokens early, even those with high confidence gaps, leads to error propagation, as these predictions may lack sufficient spatial grounding. Interestingly, the random strategy outperforms the Non-Neighbor variant, but it is still lagging behind our proposed accelerated method. This indicates that our locality-aware dynamic rescue strategy provides the optimal balance, ensuring that the accelerated decoding trajectory respects the structural dependencies of the image.
@@ 5 Conclusion and Future Work
While LADR demonstrates the potential of exploiting image spatial locality for acceleration of parallel decoding, our current method still has some limitations. The implementation relies on empirically determined hyperparameters, such as the confidence threshold $\tau$ and rescue ratios $\alpha$. These values were selected to validate the core hypothesis that spatial neighbors facilitate faster convergence rather than to locate the global optimum. Furthermore, as a training-free acceleration method, LADR’s performance is influenced by the foundational backbone. To prevent unfaithful generation, existing post-hoc alignment frameworks can be integrated to ensure prompt-image faithfulness.
@@ A.1 Proof of Theorem 1 (Margin-based Error Bound)
We aim to find the upper bound for the probability of error $P(\mathcal{E})=1-p_{(1)}$. This is strictly equivalent to finding the minimum possible value for the confidence $p_{(1)}$ subject to the given margin constraints.
@@ 1. Define Constraints: The probability distribution must sum to 1, and the margin condition must hold:
$p_{(1)}+p_{(2)}+\sum_{k=3}^{K}p_{(k)}$ $=1,$ (12)
$p_{(1)}-p_{(2)}$ $\geq\tau\implies p_{(2)}\leq p_{(1)}-\tau.$ (13)
2. Worst-Case Analysis for $p_{(1)}$: To minimize $p_{(1)}$, we must maximize the probability mass assigned to the remaining classes, particularly the closest competitor $p_{(2)}$. The most extreme case occurs when the entire probability mass is concentrated in the top two classes, meaning $\sum_{k=3}^{K}p_{(k)}\approx 0$.
Under this worst-case assumption, we have:
$p_{(1)}+p_{(2)}\approx 1.$ (14)
3. Deriving the Lower Bound of Confidence: Substituting the margin constraint from Eq. (13) into Eq. (14), we obtain:
$1$ $=p_{(1)}+p_{(2)}$
$\leq p_{(1)}+(p_{(1)}-\tau)$
$=2p_{(1)}-\tau.$ (15)
Rearranging the inequality gives the lower bound for $p_{(1)}$:
$2p_{(1)}$ $\geq 1+\tau$
@@
4. Calculating the Error Bound: The probability of error is the complement of the top-1 probability. Using Eq. (16), we bounded the error as:
$P(\mathcal{E})$ $=1-p_{(1)}$
$\leq 1-\frac{1+\tau}{2}$